# The Art of Storytelling: Evaluating AI Models for Storytelling

### v5

@@ 5.5 Final Design
Based on extensive evaluation of each component of our system, we selected the following models: for the Writer, we use Gemma-2-9b; for the Reviewer, GPT-4o; for the Narrator, XTTSv2; GPT-4o serves as both the Movie Director and Music Director; MusicGen-Large (Copet et al., 2023) as the Musician; and CogXVideo-5b, with Animated style, is used for animation generation. This configuration ensures an optimal balance of quality, performance, and safety across all stages of the storytelling process.
6 Conclusion
Our paper presents a novel multimodal multi-agent system to generate high-quality stories for school children. The initial evaluation results suggest promising results for all three forms of media generated and guide our work to a more robust and complete pipeline. Our tool aims to give an immersive educational experience to the target user base, that will not only give them an entertaining mode of learning but also maximize their pedagogical goals. By letting the users interact with our system in the next phase of our research, we will be able to crowd source evaluation data, that will help us solidify the current application, our evaluation metrics, and our long-term goals with the system.
7 Future Work
In future work, several enhancements could be made to improve the system. One key area is the integration of image inputs, where drawings provided by children could be used as prompts for story generation. Additionally, incorporating different voices for distinct character dialogues could greatly enhance the immersion of the generated audio-visual stories. Furthermore, adding AudioGen (Copet et al., 2023) as the Foley Artist agent to incorporate sound effects based on the video content would bring the stories to life.
@@ 8 Limitations
The system has several limitations that need to be addressed to enhance its performance and effectiveness. One major limitation is the need for a more comprehensive dataset to robustly test the child-appropriateness of the generated content. Additionally, the generated animations, while not inappropriate, often suffer from visual distortions and lack the smoothness and coherence required for engaging and child-friendly storytelling. These visual inconsistencies can detract from the immersive experience and may not fully align with the intended narrative quality. Another significant limitation is the high generation time for animations, with each six-second video requiring approximately three minutes to render.
@@ Appendix B System Screen
Appendix C Score Breakdown
@@ Appendix D System Prompts
Table 9 shows the system prompt for the Writer and Reviewer Agent. Table 10 shows the system prompt for Movie Director and Music Director.
Figure 6: The system’s front-end design.
@@
Agent System Prompt
@@
Agent System Prompt

### v6

@@ 5.5 User Study
Based on the evaluation of each component of our system, we selected the following models: for the Writer, we use Gemma-2-9b; for the Reviewer, GPT-4o; for the Narrator, XTTSv2; GPT-4o serves as both the Movie Director and Music Director; MusicGen-Large (Copet et al., 2023) as the Musician; and CogVideoX-5b, with Animated style, is used for animation generation. This configuration ensures an optimal balance of quality, performance, and safety across all stages of the storytelling process.
5.5.1 Children’s Responses.
Quantitative feedback from the three children is summarized below (average scores on a 1–5 scale). The quantitative results indicate that the children found the stories generally easy to understand and enjoyable, with consistently high ratings across comprehension, overall experience, and willingness to recommend the system (all averaging 4.0). The slightly lower scores for animations (3.3) and narration quality (3.7) suggest that while the multimodal elements were engaging, there is room for improvement in the visual polish and delivery of the audio components.
  1. 1.
    How easy was the story to understand? = 4.0
  2. 2.
    How likely are you to recommend this system to your friends? = 4.0
  3. 3.
    How much did you like the animations? = 3.3
  4. 4.
    How much did you like the way the story was told (e.g., narration, voice)? = 3.7
  5. 5.
    How much would you rate the generated story itself? = 4.0
  6. 6.
    Overall, how would you rate your experience with the storytelling system? = 4.0
Qualitative feedback showed that each child learned something different: one noted it was easy to “make this story by parts,” another reflected that “you shouldn’t do things behind your parents’ back,” and a third emphasized “learned new English words.” For improvements, children suggested “making the final output appear more like a book-style reading story,” “adding humor to prevent boredom,” and “improving both animation quality and narration expressiveness.” Overall, their responses highlight the system’s potential for engaging, age-appropriate learning while pointing to areas for refinement in visuals and narration
5.5.2 Parents’ Responses.
Quantitative feedback from the three parents is summarized below (average scores on a 1–5 scale). The results show that parents considered the content highly appropriate for their children (5.0) and were generally satisfied with both their child’s engagement (4.3) and the overall system experience (4.3). Willingness to recommend the system to other parents (4.0) and satisfaction with the design (3.7) received slightly lower scores, suggesting that while parents valued the educational content and engagement, there remains room for improvement in interface design and usability.
  1. 1.
    How appropriate was the content for your child’s age? = 5.0
  2. 2.
    How engaged was your child while using the system? = 4.3
  3. 3.
    How likely are you to recommend this system to other parents? = 4.0
  4. 4.
    How much did you like the system’s design (e.g., visuals, ease of use)? = 3.7
  5. 5.
    Overall, how satisfied are you with the storytelling system? = 4.3
Qualitative feedback showed that children generally reacted positively to the system, though enthusiasm varied. One parent noted their child was “excited,” another that she “enjoyed it but not as much as I thought she would, especially given she loves to write stories,” while a third observed he “was engaged and was looking forward to what will be generated.” Parents also highlighted learning benefits such as navigating new digital tools, adding structure to story writing, and gaining new English vocabulary. They agreed the system could enhance writing skills, vocabulary, and pronunciation through narration. Overall, parents valued the system’s educational potential and meaningful engagement, while pointing to design and presentation as areas for improvement. These findings suggest that the system may support and encourage learning motivation, though direct educational outcomes were not measured in the present study.
6 Conclusion & Future Work
We presented a multimodal multi-agent system for generating high-quality stories for school children. Our evaluation results suggest strong potential across all media types and highlight steps toward a more robust pipeline. The tool provides an immersive educational experience that is both entertaining and pedagogically valuable. Future work includes integrating image inputs (e.g., children’s drawings) as prompts, adding distinct voices for character dialogues to enhance immersion, and incorporating AudioGen (Copet et al., 2023) as a Foley Artist to generate context-appropriate sound effects behind the animations. In addition, improving the quality of the TTS and TTV modules, as highlighted in the user study, remains a key goal for enhancing the overall storytelling experience.
Beyond the application itself, our study provides a comparative qualitative evaluation of multiple generative models across three media modalities (text, visual, and audio), highlighting their relative strengths, weaknesses, and capability differences. This evaluation framework can serve as a reference benchmark for assessing future multimodal storytelling systems and related generative models. Moreover, the human-in-the-loop storytelling process explored in this work provides a foundation for future research on AI-assisted data collection and human-AI co-creative educational systems.
@@ 7 Limitations
The system has several limitations that need to be addressed to enhance its performance and effectiveness. One major limitation is the need for a more comprehensive dataset to robustly test the child-appropriateness of the generated content. Additionally, the generated animations, while not inappropriate, often suffer from visual distortions and lack the smoothness and coherence required for engaging and child-friendly storytelling. These visual inconsistencies can detract from the immersive experience and may not fully align with the intended narrative quality. In addition, the creativity of the generated models is also limited by the training dataset. This may be improved by using a wider genre of stories rather than a large dataset with many similar stories. The fairytales and folktales from Project Gutenberg that were used to train the Writer LLM share many similar themes and plots. This necessitates the creation of a newer, diverse dataset of stories. Another significant limitation is the high generation time for animations, with each six-second video requiring approximately three minutes to render. In our study, the practical impact of this limitation was mitigated for younger participants (ages 6-9). The combination of multiple media elements (text, audio, and visual) increased their content-consumption time, giving animations buffer time to render. However, this is unlikely to generalize to older participants with faster reading speeds. Responsiveness could be significantly improved by reducing video duration or replacing it entirely with static images, replicating a picture-book experience. Evaluating these alternatives constitutes an important direction for future work.
@@ Appendix B System Screen
Appendix C Score Breakdowns
@@ Appendix D System Prompts
Table 11 shows the system prompt for the Writer and Reviewer. Table 12 shows the system prompt for Movie Director and Music Director.
@@
Model System Prompt
@@
Model System Prompt